OpenStudy (anonymous):

Topic: \(\textbf{Calculus 2}\) Evaluate the indefinite integral as an infinite series. \[\int \frac{\cos x − 1}{x} \ \ \ dx \] What the heck? I mean putting these in Taylor series forms are hard enough, but this? Not even sure where to start here >_< It worries me because the final is coming up and this is a hole in my knowledge... So, um, help?

OpenStudy (anonymous):

Well, could you start with the infinite series for cosine x?

OpenStudy (anonymous):

You'd have (1/x) (-x^2/2! + x^4/4! .....)

OpenStudy (anonymous):

\[ \int \frac{\cos x}{x}dx - \int \frac{1}{x}dx \]

OpenStudy (anonymous):

Let's see... \[\sum_{n=0}^{\infty} (-1)^n \frac{x^{2n+1}}{(2n+1)!}, \text{for all x = } \mathbb{R}\]

OpenStudy (anonymous):

That's the MacLaurin equivalent for cosine

OpenStudy (anonymous):

\[ \frac{ \cos x}{x} = \frac{1}{x} \sum_{n=0}^\infty (-1)^n \frac{x^{2n}}{(2n)!} \]

OpenStudy (anonymous):

Is it? Did I look at the wrong one? (-:

OpenStudy (anonymous):

oh yes that's the hyperbolic one! SOrry!

OpenStudy (anonymous):

Oh no wait I have the wrong one, I wrote sine by mistake

OpenStudy (anonymous):

Arg soo confusing lol

OpenStudy (anonymous):

lets both check again then hehe.

OpenStudy (anonymous):

yes I believe I have the right one

OpenStudy (anonymous):

I thought you couldn't just do 1/x , you'd have to convert that to a series as well, yes?

OpenStudy (experimentx):

\[ \cos x = \sum_{n = 0}^\infty {(-1)^nx^{2n}\over (2n)!}\] which makes \[ \int \sum_{n = 1}^\infty {(-1)^nx^{2n-1}\over (2n)!} dx = {(-1)^nx^{2n}\over 2n(2n)!}\]

OpenStudy (experimentx):

+c

OpenStudy (anonymous):

Anti-derivative power rule?

OpenStudy (anonymous):

\[ \frac{ \cos x}{x} = \frac{1}{x} \sum_{n=0}^\infty (-1)^n \frac{x^{2n}}{(2n)!} \] Distribute it out \[ \frac{\cos x}{x} = \sum_{n=0}^\infty \frac{(-1)^nx^{2n-1}}{(2n)!} \]

OpenStudy (anonymous):

yes that's it @agentx5

OpenStudy (anonymous):

all has mistake!

OpenStudy (anonymous):

where @mahmit2012 ?

OpenStudy (anonymous):

all !but it is time to pray i wll come back and show you

OpenStudy (experimentx):

\[ \int \sum_{n = 1}^\infty {(-1)^nx^{2n-1}\over (2n)!} dx = \sum_{n=1}^\infty {(-1)^nx^{2n}\over 2n(2n)!} \]

OpenStudy (anonymous):

after pray !

OpenStudy (experimentx):

lol ... where did i make mistake??

OpenStudy (anonymous):

Did we do the same @experimentX ? besides that you completed it already.

OpenStudy (anonymous):

I think we did.

OpenStudy (experimentx):

n=1, n=0

OpenStudy (anonymous):

yes but obviously "ALL" is wrong :D

OpenStudy (anonymous):

first of all the original question has interval [0,inf) and it has many different solution to get the answer. and if you are looking for undefined integral so just a matching series is the answer ! but with a log function.

OpenStudy (anonymous):